If an amplifier working with ±5 V DC power supplies requires a minimum output voltage of 2.3 V across a load resistance of 185 ohms, calculate the minimum average current to be drawn from each power supply, if the efficiency of the amplifier is 50 %.


1
Expert's answer
2021-05-26T15:34:02-0400

Efficiency , "\eta=50%" %"=0.50"

P"_{out}=\frac{V^2_{out}}{R}=\frac{2.3^2}{185}=0.02859"

"\eta=\frac{P_{out}}{P_{dc}}=0.50=\frac{0.02859}{P_{dc}}"

"P_{dc}=\frac{0.02859}{0.5}=0.057189"

"V_{cc}I_{c}=P_{dc}"

"I_c=\frac{P_{dc}}{V{cc}}=\frac{0.057189}{5}=0.0114378A"